If f(x) = 4x - 3,
find f(-1)
findf(a-2)

Answer by rmromero(383)


Solution:
From the equation, f(x) = 4x - 3, Find f(-1). This shows that your x = -1 so what you are going to do is substitute -1 to x. so we will have
f(-1) = 4(-1) -3
= -4 -3
= -7
therefore your f(x) = -7 if f(-1).
what if you have f(a-2), This shows that x = a-2 so substitute a-2 to x from the equation f(x) = 4x - 3
f(a-2) = 4(a-2) - 3 Distribution property
= 4a - 8 - 3 Addition property
= 4a - 11
therefore f(x) = 4a - 11 if f(a-2)